# SeFFeC: Semantic Facial Feature Control for Fine-grained Face Editing

### v1

@@ Abstract
We propose Semantic Facial Feature Control (SeFFeC) – a novel method for fine-grained face shape editing. Our method enables the manipulation of human-understandable, semantic face features, such as nose length or mouth width, which are defined by different groups of facial landmarks. In contrast to existing methods, the use of facial landmarks enables precise measurement of the facial features, which then enables training SeFFeC without any manually annotated labels. SeFFeC consists of a transformer-based encoder network that takes a latent vector of a pre-trained generative model and a facial feature embedding as input, and learns to modify the latent vector to perform the desired face edit operation. To ensure that the desired feature measurement is changed towards the target value without altering uncorrelated features, we introduced a novel semantic face feature loss. Qualitative and quantitative results show that SeFFeC enables precise and fine-grained control of 23 facial features, some of which could not previously be controlled by other methods, without requiring manual annotations. Unlike existing methods, SeFFeC also provides deterministic control over the exact values of the facial features and more localised and disentangled face edits.
Keywords:
Face Editing Facial Landmarks Deterministic Editing
Figure 1: We propose SeFFeC – a novel method for fine-grained control over a large variety of semantic face shape features, such as eye width or chin shape, without requiring manual labels or user effort. In contrast to existing methods, SeFFeC can precisely control the desired degree of change.
@@ 1 Introduction
Controllable face generation is an active area of research with broad applications in digital image editing. Recent methods, such as StyleGAN [23, 24], have been shown to learn disentangled latent image spaces that facilitate controllable image generation [8, 41, 50, 55] and, because of this, there has been a growing interest in methods that enable users to control the generative process. There are a number of different classes of methods, from unsupervised that automatically discover disentangled dimensions and control specific parts of the face [16, 43, 34] to mask-based methods that use segmentation masks to control the face geometry [13, 44, 29, 30, 46, 47], text-based methods that use natural language [56, 22, 36, 20, 45, 3], 3D-based methods that translate a 3D face model to a real image [50, 10, 49, 33, 27], or attribute-based approaches [7, 57, 32, 12, 42, 60, 1, 55] that can control specific appearance attributes. While all these methods can produce impressive results, they have several limitations that prevent users from easily and controllably changing the appearance of a face. Attribute- and mask-based methods are tedious as they require either manual annotations or the skill to manipulate segmentation masks. While unsupervised methods overcome these limitations, they are often model-dependent and dimensions to manipulate the desired attributes might not be discovered or are still entangled with other attributes. 3D-based methods are great for novel-view synthesis, lighting manipulation or transferring expressions, but are not capable to perform semantic face shape editing. Moreover, none of these methods allow users to observe and edit precise values of the facial features they want to edit (e.g. the width of the eyes), which prevents deterministic semantic control.
@@
To address these limitations we propose Semantic Facial Feature Control (SeFFeC). Our method can manipulate 23 semantic face features that cover key characteristics of human face geometry, such as the width and openness of the eyes and the mouth, the gaze direction, or the chin and eyebrow shape (see Figure 1). Similarly to action units from the Facial Action Coding System, which describes how the face moves when embodying different emotions [11], our proposed semantic face features are derived from and describe the relation between facial landmarks and face characteristics. Unlike existing methods, given that our novel semantic face features are derived from landmarks, we are able to measure the exact values, which allows us to train SeFFeC without any manual annotations. We inject a Transformer encoder network into a pre-trained StyleGAN2 model that predicts a manipulation vector based on a semantic face feature embedding and the latent code of an image. This vector is subsequently scaled and added to the original latent code to change a specific face feature by a desired amount. To train the model we propose a novel semantic face feature loss that enforces the change in the desired face feature. Furthermore, we estimated the correlations of our semantic face features on real human faces to dynamically regularise SeFFeC  allowing it to modify naturally correlated measures jointly.
@@
We report qualitative and quantitative results showing that SeFFeC allows for deterministic manipulation of face features with a high degree of disentanglement. We compared to two state-of-the-art baselines and show that SeFFeC generates more precise and localised face editing results. In addition, we developed a simple user interface that consists of multiple sliders, similar to common digital character editing tools [40, 2] that provide users with an easy-to-use interface with precise control over the face feature values (see video demonstration: https://youtu.be/1xQb9zDs-Kg - anonymised for review), paving the way for more sophisticated digital face editing tools than possible so far [^1] Project code, models, as well as an interactive demo will be released upon acceptance..
@@ 2 Related Work
Most unsupervised methods aim at decomposing the latent space [16, 34, 6] or weights [43] of a generative model to identify semantic editing dimensions. While such methods do not require labelled data, only a limited amount of meaningful semantic directions can be discovered and the level of entanglement is typically higher compared to supervised methods. DragGAN [35] allows users to edit the shape of a face by moving facial landmarks. While this allows for a variety of shape edits, this requires significant user effort, especially for edits involving multiple landmarks. Moreover, DragGAN cannot be used for automated face editing. While the current landmarks can be automatically detected, the target location of the landmarks are unknown and to be determined by the users. Instead, we directly specify semantic features using landmarks, allowing for effortless user control and precise automatic manipulation.
Mask-based Face Editing.
@@
Mask-based methods condition a generative model on face masks to provide fine-grained control. One common approach is to allow users to erase parts of the face and sketch the desired outlines of facial features. Subsequently, a generative model performs image inpainting based on the sketch mask [37, 5, 4]. Another approach is to condition on face segmentations, where facial features like the eyes and mouth are separated by masks. This allows users to edit faces by manipulating the face segmentation masks [13, 44, 29, 30, 46, 47]. Mask-based methods require significant user effort and skill to manipulate the masks to perform the desired edits. With our method a simple slider interface can be used for face editing, which does not require any special skill.
Text-based Face Editing.
@@
Xia et al. [56] and Huang et al. [22] combined the ideas from mask-based techniques with text-guided image edits. Patashnik et al. [36] combined a pre-trained generator and a powerful CLIP [39] image-language encoder for improved text-guided editing capabilities. Subsequent work improved upon this to perform more localised edits [20] or to generalise to more diverse text prompts [45]. Text-based face editing methods provide a simple language interface to facilitate a large variety of editing possibilities. However, these methods do not allow for very precise and fine-grained edits. Instead, we define a large set of semantic face features that provides users with very fine-grained and precise control over the face shape.
Attribute-based Face Editing.
@@
Early work on attribute-based face editing focused on binary control of attributes defined during inference [7, 57, 32, 62, 14, 58], e.g. open or closed mouth. Because many of these methods inflicted unwanted global changes to the faces, subsequent works used attention mechanisms to learn which face regions have to be modified for an attribute change, resulting in more disentangled and localised edits [62, 28, 53, 17]. Other works used attribute classifiers during training for improved editing results [18, 48, 12]. Generative models conditioned on the attributes have typically limited progressive editing control. Therefore, Shen et al. [42] trained SVMs to separate face attributes in the generator’s latent space with the SVM normal vectors providing global semantic directions along which the corresponding attributes could be manipulated smoothly. Han et al. [15] improved these global directions with a local iterative search based on the current image.
@@
Another approach is to learn how to modify the latent vectors of a pre-trained generative face model to perform semantically plausible face edits [60, 21, 25, 1, 55]. Yang et al. [59] proposed a method that did not require binary attributes but sets of images that each shared a particular attribute. While significant progress has been made in attribute-based face editing, two main disadvantages remain: First, latest methods can only control attributes for which enough manually labelled data exists. Manual labelling at scale is not only tedious, time-consuming, and costly, it can also be challenging to provide binary labels for features like chin shape for which no clear boundary exists. Second, while current methods allow for progressive control over attributes, it remains unclear for users how far to move along a semantic dimension to achieve a desired result, e.g. to fully close the mouth. Our method addresses both limitations as it does not require any manually labelled data and provides semantically meaningful, deterministic control over the face shape feature values.
@@ 3 Method
The core idea of our Semantic Facial Feature Control (SeFFeC) method is to allow to control face characteristics through a set of semantic face features in a deterministic way. In the following, we first define our semantic face features and then present how SeFFeC is able to manipulate the shape of a face based on these features.
@@ 3.1 Landmark-based Face Features
Table 1: List of the proposed semantic face features based on different facial landmarks. The characters $x$ and $y$ indicate if the x- or y-coordinate of a particular landmark is used, while the subscripts refer to the landmarks shown in Figure 2 (right).
@@
Eyebrow thickness $(y_{41}-y_{34})+(y_{38}-y_{37})+(y_{50}-y_{42})+(y_{47}-y_{45})\phantom{+}$
Eyebrow shape $(y_{33}-y_{35})+(y_{37}-y_{35})+(y_{42}-y_{44})+(y_{46}-y_{44})\phantom{+}$
@@
We define semantic face features $\mathcal{M}$ based on commonly used 98 landmarks available in the Wider Facial Landmarks in the Wild (WFLW) dataset [54] (see Figure 2 right). Inspired by common digital face creation tools [40, 2] that provide an interface with multiple sliders to manipulate semantic face features, we define a total of 23 features $m\in\mathcal{M}$ (see Table 1). $x$ and $y$ indicate which dimension of a landmark we used for calculating the feature while the subscripts refer to the corresponding landmark as shown in Figure 2 (right). The features fall into three different categories: absolute distance, relative distance, and relative anchor distance. Most features are defined based on the relative distance between two or more landmarks, such as eye width or nose length. The advantage of relative distance features is that they are invariant to head translations. Given a lack of clear reference landmarks, we additionally define five absolute distance features that encode the absolute position of facial landmarks in the face image: horizontal and vertical pupil position, eyebrow height, mouth height, and chin length. The disadvantage of these features is that they are not translation-invariant. For example, translating the whole face in an image downwards also increases the chin length. As we will show later, however, these features are still effective for controlling the desired facial features. While it is possible to define relative reference landmarks for these features, e.g. by relating the landmark used for chin length ($y_{16}$) with the position of the mouth, this can lead to undesired side effects because that feature would be affected by both chin length and mouth position. For eyebrow shape, chin shape, and mouth shape we define translation-invariant features that relate two landmarks to a third one and, as such, allow us to control the angle between them and the shape of the underlying facial feature. Any differentiable function to combine landmarks can be used, potentially allowing for many more interesting semantic face features.
@@
Figure 2: Overview of our method. Shown in grey is the StyleGAN2 architecture. We inject a transformer encoder network between the StyleGAN mapping and synthesis network (green) that can modify the latent vector $w^{+}$ based on the desired face feature embedding $e_{j}$ by adding a semantic manipulation vector $s_{e}$. This manipulation vector is scaled by $k$, a scalar predicted by the scaling network based on the current and target face feature values $m_{j}$ and $m_{j}^{t}$. The landmark detector and calculated face features (blue) are only required during training of the components highlighted in green.
3.2 Semantic Face Feature Control
@@
Our goal is to develop a method that allows to manipulate the latent vector of an image $I$ in such a way that only a desired semantic face feature is changed towards a target value, thus generating a new image $I_{\text{edit}}$. Figure 2 shows the overall architecture of the proposed method to achieve this goal. The grey components are from the state-of-the-art generative network StyleGAN2 [24] that was pre-trained on the Flickr-Faces-HQ dataset (FFHQ) [23] to generate high-quality images of human faces. A mapping network first maps latent vectors $z\in\mathbb{R}^{512}$ drawn from a standard normal distribution, $z\sim\mathcal{N}(0,1)$, into $\mathcal{W}$ space. The vector $w\in\mathcal{W}$ is then repeated 18 times, $w^{+}\in\mathcal{W}^{+}\in\mathbb{R}^{512\times 18}$ and input to the generative network that subsequently generates the image. Prior work has shown that the $\mathcal{W}^{+}$ space is highly disentangled and can be modulated to perform various semantic image manipulations [8, 41, 50, 55].
@@
such that $w^{+}_{\text{edit}}$ translates to the same image as $w^{+}$ except that the value of the selected semantic face feature $m$ is changed to the target value. For this, we input $w^{+}$ as a sequence of 18 $w$ vectors with an additional embedding vector $e_{j}\in\mathbb{R}^{512}$ into the Transformer. The vector $e_{j}$ represents an embedding that encodes information about the semantic face feature $m_{j}\in\mathcal{M}$ to be manipulated. The embeddings are produced by an additional embedding layer that is trained end-to-end and encodes the 23-dimensional one-hot face feature vectors based on our semantic features defined in Table 1. The Transformer consists of multiple layers of multi-head self-attention to learn how to manipulate $w^{+}$ based on $e_{j}$ to generate $s_{e}$. The semantic manipulation vector $s_{e}$ can be extracted from the Transformer’s output by ignoring the last element in the output sequence, which belongs to the input embedding vector. A scaling network takes the embedding vector as well as the current and target values for the face feature we want to change, $m_{j}$ and $m_{j}^{t}$, as input. It predicts a scaling factor $k$ which is multiplied with $s_{e}$ as defined in Equation 1, allowing us to deterministically change the face feature value by a desired amount.
During training of the scaling network, embedding layer and Transformer encoder, our method requires a differentiable landmark detector to be able to calculate the face features $M$ (highlighted in blue in Figure 2). The weights of StyleGAN2 as well as the landmark detector model are frozen and only used to calculate the gradient to update the weights of the components highlighted in green. We define the full loss function $\mathcal{L}$ to train the network as follows:
@@
with $\lambda_{\text{pix}}$, $\lambda_{\text{feat}}$ and $\lambda_{\text{SFF}}$ representing scalars to weight the different loss terms. $\mathcal{L}_{\text{pix}}$ is defined as the pixel based mean squared error (MSE) between the original image $I$ and the modified image $I_{\text{edit}}$ which aims at preserving the original image as much as possible. This loss is zero if the Transformer predicts $s_{e}=\overrightarrow{0}$, i.e., the network should perform as few changes as possible on $w^{+}$ to fulfil the other constraints. $\mathcal{L}_{\text{feat}}$ is defined as the feature based MSE between $I$ and $I_{\text{edit}}$, where we consider the features of the last style block in the StyleGAN2 generator. Similarly to $\mathcal{L}_{\text{pix}}$, this loss is zero for $s_{e}=\overrightarrow{0}$ and forces the network to make minimal changes. The feature supervision is less rigid than the pixel based loss and allows the model to be more flexible in changing the face, while still preserving the appearance of the original face. $\mathcal{L}_{\text{SFF}}$ is the semantic face feature loss and responsible for the network to learn how to modify the semantic face features. It is defined as:
$\mathcal{L}_{\text{SFF}}=\text{ MSE}(m_{j}^{\text{p}},m_{j}^{\text{t}})+\lambda_{\text{reg}}\sum_{i=1,i\neq j}^{23}\text{MSE}(m_{i}^{\text{p}},m_{i}^{\text{t}})*(1-\lambda_{\text{cor}}\ *\ |c(m_{i},m_{j})|).$ (3)
The first term calculates the MSE between the face feature $m_{j}^{p}$ of the edited face $I_{\text{edit}}$ and the target value for the feature $m_{j}^{t}$. To calculate $m_{j}^{p}$, the generated image is passed through a differentiable landmark detector in the forward-pass as illustrated in Figure 2 to predict the landmarks for $I_{\text{edit}}$. While this term encourages the model to change $w^{+}$ to manipulate the feature $m_{j}$, we need to ensure that all other face features remain unchanged. We therefore add a regularising term with weighting $\lambda_{\text{reg}}$ that sums the MSE between predicted and target values for all other features where $m_{i}^{t}$ is fixed to the value of the original image $I$. Furthermore, the regularising term is multiplied by the inverse of the absolute Person correlation coefficient $c(m_{i},m_{j})$ between the feature to be changed $m_{j}$ and those to be kept fixed $m_{i}$. Given that not all of our semantic features are necessarily fully disentangled given human anatomy, we estimate the correlations between the features on the FFHQ dataset [24] and introduce scalar $\lambda_{\text{cor}}$ that controls how much the correlation can relax the regularising term: If two features are strongly correlated the regularising term between them is relaxed and the network is less penalised for changing the correlated features $m_{i}$ together with the target feature $m_{j}$.

### v3

# UP-FacE: User-predictable Fine-grained Face Shape Editing

@@ Abstract
We present User-predictable Face Editing (UP-FacE) – a novel method for predictable face shape editing. In stark contrast to existing methods for face editing using trial and error, edits with UP-FacE are predictable by the human user. That is, users can control the desired degree of change precisely and deterministically and know upfront the amount of change required to achieve a certain editing result. Our method leverages facial landmarks to precisely measure facial feature values, facilitating the training of UP-FacE without manually annotated attribute labels. At the core of UP-FacE is a transformer-based network that takes as input a latent vector from a pre-trained generative model and a facial feature embedding, and predicts a suitable manipulation vector. To enable user-predictable editing, a scaling layer adjusts the manipulation vector to achieve the precise desired degree of change. To ensure that the desired feature is manipulated towards the target value without altering uncorrelated features, we further introduce a novel semantic face feature loss. Qualitative and quantitative results demonstrate that UP-FacE enables precise and fine-grained control over 23 face shape features.
Figure 1: UP-FacE  allows for fine-grained control over various face shape features, such as eye width, eye and mouth openness, or eyebrow thickness. Qualitative and quantitative results demonstrate that UP-FacE enables precise and fine-grained control over 23 face shape features. In stark contrast to existing methods that require trial and error editing of face features, edits with UP-FacE are predictable by the human user. That is, users can control the desired degree of change precisely and deterministically and know upfront the amount of change required to achieve a certain editing result. In addition, our method enables both isolated progressive (i.e. of the same feature) and sequential (i.e. multiple different features) edits without altering other (unrelated) facial features.
@@ 1 Introduction
Computational face editing is an active area of research with broad applications, such as digital image editing. Powerful generative models such as StyleGAN [22, 23] can generate high-quality face images and have been shown to learn disentangled latent image spaces that facilitate controlled face editing [8, 40, 49, 54]. Because of this, there has been a growing interest in methods that enable users to control the generative process. Various methods have been proposed for controllable face editing, from unsupervised ones that automatically discover disentangled latent dimensions to control specific parts of the face [15, 42, 33] over mask-based methods that use segmentation masks to control the face geometry [13, 43, 28, 29, 45, 46], to text-based methods that use natural language [55, 21, 35, 19, 44, 3], 3D-based methods that translate a 3D face model to a real image [49, 10, 48, 32, 26], or attribute-based approaches [7, 56, 31, 12, 41, 59, 1, 54] that can control specific appearance attributes.
While all of these methods can produce impressive results, they also have fundamental limitations that prevent users from easily and deterministically editing the face shape: Attribute- and mask-based methods are tedious as they require either manual annotations or sufficient skill to manipulate segmentation masks. While unsupervised methods overcome these limitations, they are often model-dependent, and dimensions to manipulate the desired attributes might not be discovered or might be entangled with other attributes. 3D-based methods are effective for novel-view synthesis, lighting manipulation, and transferring expressions but require significant 3D modelling efforts for face shape editing. Most importantly, none of these methods permit users to know the outcome of a particular face edit upfront. Instead, users must manipulate facial features via trial and error until they are satisfied with the result, resulting in a tedious and error-prone face editing process.
@@
To address these limitations, we present User-predictable Face Editing (UP-FacE). UP-FacE can manipulate 23 face features that describe key characteristics of human face geometry, such as the eye and mouth width and openness, or the chin and eyebrow shape (see Figure 1). Similar to action units from the Facial Action Coding System, which describe how face appearance changes for different emotions [11], we introduce semantic face features that are derived from and describe the relation between facial landmarks and face characteristics. Unlike existing methods, these semantic features are derived from landmarks, enabling us to quantify the exact facial feature values. This ability allows us to train UP-FacE without the need for any manual attribute annotations. To train our model, we propose a novel semantic face feature loss that enforces the change in the desired face feature. Furthermore, we estimated the correlations of the semantic face features on real human faces to dynamically regularise UP-FacE  allowing it to modify naturally correlated features jointly.
@@
We report qualitative and quantitative results showing that UP-FacE allows for predictable manipulation of facial features with a high degree of disentanglement. Most crucially and novel compared to previous methods, our capability to measure and set precise facial feature values enables predictable user editing outcomes: Users can control the desired degree of change precisely and deterministically and know upfront the amount of change required to achieve a certain editing result. In addition, we developed a user interface that consists of multiple sliders, similar to common digital character editing tools [39, 2] that provide users with an easy-to-use interface with precise control over the face feature values (see video demonstration: https://youtu.be/xSXAJP1M3ew), paving the way for more user-friendly digital face editing tools.
@@ 2 Related Work
Unsupervised Face Editing. Most unsupervised methods aim at decomposing the latent space [15, 33, 6] or weights [42] of a generative model to identify semantic editing dimensions. While such methods do not require labelled data, only a limited amount of meaningful semantic directions can be discovered and the level of entanglement is typically higher compared to supervised methods. DragGAN [34] allows users to edit the shape of a face by moving facial landmarks. While this allows for a variety of shape edits, this requires significant user effort, especially for edits involving multiple landmarks. Instead, we specify semantic features using landmarks, which enables UP-FacE to learn how to move these landmarks, thereby facilitating effortless user control.
@@
Mask-based Face Editing. Mask-based methods condition a generative model on face masks to provide fine-grained control. One common approach is to allow users to erase parts of the face and sketch the desired outlines of facial features. Subsequently, a generative model performs image inpainting based on the sketch mask [36, 5, 4]. Another approach is to condition face segmentations, where facial features like the eyes and mouth are separated by masks. This allows users to edit faces by manipulating the face segmentation masks [13, 43, 28, 29, 45, 46]. Mask-based methods require significant user effort and skill to manipulate the masks to perform the desired edits. With our method, a slider interface can be used for face editing, which does not require any special skill.
@@
Text-based Face Editing. Xia et al. [55] and Huang et al. [21] combined the ideas from mask-based techniques with text-guided image edits. Patashnik et al. [35] combined a pre-trained generator and a powerful CLIP [38] image-language encoder for improved text-guided editing capabilities. Subsequent work improved upon this to perform more localised edits [19] or to generalise to more diverse text prompts [44]. Text-based face editing methods provide a simple language interface to facilitate a large variety of editing possibilities. However, these methods do not allow for very precise and fine-grained edits.
@@
Attribute-based Face Editing. Early work on attribute-based face editing focused on binary control of attributes defined during inference [7, 56, 31, 61, 14, 57, 61, 27, 52, 16], e.g. open or closed mouth. Other works used attribute classifiers during training for improved editing results [17, 47, 12]. Generative models conditioned on the attributes have typically limited progressive editing control. Therefore, Shen et al. [41] trained SVMs to separate face attributes in the generator’s latent space with the SVM normal vectors providing global semantic directions along which the corresponding attributes could be manipulated smoothly.
@@
Another approach is to learn how to modify the latent vectors of a pre-trained generative face model to perform semantically plausible face edits [59, 20, 24, 1, 54]. Yang et al. [58] proposed a method that did not require binary attributes but sets of images that each shared a particular attribute. While significant progress has been made in attribute-based face editing, two main disadvantages remain: First, the latest methods can only control attributes for which enough manually labelled data exists. Manual labelling at scale is not only tedious, time-consuming, and costly, but it can also be challenging to provide binary labels for features like chin shape for which no clear boundary exists. Second, while current methods allow for progressive control over attributes, it remains unclear for users how far to move along a semantic dimension to achieve a desired result, e.g. to fully close the mouth. Our method addresses both limitations as it does not require any manually labelled data and provides semantically meaningful, user-predictable control over the face shape features.
@@ 3 Method
The core idea of our User-predictable Face Editing (UP-FacE) method is to allow the control of face characteristics through a set of semantic face features in a deterministic way. In the following, we first define our semantic face features and then present how UP-FacE can manipulate the shape of a face based on these features.
@@ 3.1 Landmark-based Face Features
Eyebrow thickness $(y_{41}-y_{34})+(y_{38}-y_{37})+$
$(y_{50}-y_{42})+(y_{47}-y_{45})\phantom{+}$
Eyebrow shape $(y_{33}-y_{35})+(y_{37}-y_{35})+$
$(y_{42}-y_{44})+(y_{46}-y_{44})\phantom{+}$
@@
Table 1: List of the proposed semantic face features based on different facial landmarks. The characters $x$ and $y$ indicate if a particular landmark’s x- or y-coordinate is used, while the subscripts refer to the landmarks shown in Figure 2 (right).
@@
We define semantic face features $\mathcal{M}$ based on commonly used 98 2D landmarks available in the Wider Facial Landmarks in the Wild (WFLW) dataset [53] (see Figure 2 right). Inspired by common digital face creation tools [39, 2] that provide an interface with multiple sliders to manipulate semantic face features, we define a total of 23 features $m\in\mathcal{M}$ (see Table 1). $x$ and $y$ indicate which dimension of a landmark we used for calculating the feature, while the subscripts refer to the corresponding landmark as shown in Figure 2 (right). The features fall into three different categories: absolute distance, relative distance, and relative anchor distance. Most features are defined based on the relative distance between two or more landmarks, such as eye width or nose length. The advantage of relative distance features is that they are invariant to head translations. Given a lack of clear reference landmarks, we additionally define five absolute distance features that encode the absolute position of facial landmarks in the face image: horizontal and vertical pupil position, eyebrow height, mouth height, and chin length. The disadvantage of these features is that they are not translation-invariant. For example, translating the whole face in an image downwards also increases the chin length. However, as we will show later, these features are still effective for controlling the desired facial features. While it is possible to define relative reference landmarks for these features, e.g. by relating the landmark used for chin length ($y_{16}$) with the position of the mouth, this can lead to undesired side effects because that feature would be affected by both chin length and mouth position. For eyebrow shape, chin shape, and mouth shape, we define translation-invariant features that relate two landmarks to a third one (anchor) and, as such, allow us to control the angle between them and the shape of the underlying facial feature. Any differentiable function to combine landmarks can be used, potentially allowing for many more interesting semantic face features.
@@
Figure 2: Overview of our method. Shown in grey is the StyleGAN2 architecture. We inject a transformer encoder network between the StyleGAN mapping and synthesis network (green) that can modify the latent vector $w^{+}$ based on the desired face feature embedding $e_{j}$ by adding a semantic manipulation vector $s_{e}$. This manipulation vector is scaled by $k$, a scalar predicted by the scaling network based on the current and target face feature values $m_{j}$ and $m_{j}^{t}$. The landmark detector and calculated face features (blue) are only required during the training of the components highlighted in green.
3.2 User-predictable Face Editing
@@
Our goal is to develop a method that allows the predictable manipulation of the latent vector of an image $I$ in such a way that only a desired semantic face feature is changed towards a target value, thus generating a new image $I_{\text{edit}}$. Figure 2 shows the overall architecture of the proposed method to achieve this goal. The grey components are from the state-of-the-art generative network StyleGAN2 [23] that was pre-trained on the Flickr-Faces-HQ dataset (FFHQ) [22] to generate high-quality images of human faces. A mapping network first maps latent vectors $z\in\mathbb{R}^{512}$ drawn from a standard normal distribution, $z\sim\mathcal{N}(0,1)$, into $\mathcal{W}$ space. The vector $w\in\mathcal{W}$ is then repeated 18 times, $w^{+}\in\mathcal{W}^{+}\in\mathbb{R}^{512\times 18}$ and input to the generative network that subsequently generates the image. Prior work has shown that the $\mathcal{W}^{+}$ space is highly disentangled and can be modulated to perform various semantic image manipulations [8, 40, 49, 54].
@@
such that $w^{+}_{\text{edit}}$ translates to the same image as $w^{+}$ except that the value of the selected semantic face feature $m$ is changed to the target value. We input $w^{+}$ as a sequence of 18 $w$ vectors with an additional embedding vector $e_{j}\in\mathbb{R}^{512}$ into the Transformer. The vector $e_{j}$ represents an embedding that encodes information about the semantic face feature $m_{j}\in\mathcal{M}$ to be manipulated. The embeddings are produced by an additional embedding layer that is trained end-to-end and encodes the 23-dimensional one-hot face feature vectors based on our semantic features defined in Table 1. The Transformer consists of multiple layers of multi-head self-attention to learn how to manipulate $w^{+}$ based on $e_{j}$ to generate $s_{e}$. The semantic manipulation vector $s_{e}$ can be extracted from the Transformer’s output by ignoring the last element in the output sequence, which belongs to the input embedding vector. A scaling network takes the embedding vector and the current and target values for the face feature we want to change, $m_{j}$ and $m_{j}^{t}$, as input. It predicts a scaling factor $k$ multiplied with $s_{e}$ as defined in Equation 1, allowing us to change the face feature value by a desired amount deterministically.
During the training of the scaling network, embedding layer and Transformer encoder, our method requires a differentiable landmark detector to be able to calculate the face features $M$ (highlighted in blue in Figure 2). The weights of StyleGAN2 and the landmark detector model are frozen and only used to calculate the gradient to update the weights of the components highlighted in green. We define the full loss function $\mathcal{L}$ to train the network as follows:
@@
with $\lambda_{\text{pix}}$, $\lambda_{\text{feat}}$ and $\lambda_{\text{SFF}}$ representing scalars to weight the different loss terms. $\mathcal{L}_{\text{pix}}$ is defined as the pixel-based mean squared error (MSE) between the original image $I$ and the modified image $I_{\text{edit}}$ which aims at preserving the original image as much as possible. This loss is zero if the Transformer predicts $s_{e}=\overrightarrow{0}$, i.e., the network should perform as few changes as possible on $w^{+}$ to fulfil the other constraints. $\mathcal{L}_{\text{feat}}$ is defined as the feature-based MSE between $I$ and $I_{\text{edit}}$, where we consider the features of the last style block in the StyleGAN2 generator. Similarly to $\mathcal{L}_{\text{pix}}$, this loss is zero for $s_{e}=\overrightarrow{0}$ and forces the network to make minimal changes. The feature supervision is less rigid than the pixel-based loss and allows the model to be more flexible in changing the face while preserving the original face’s appearance. $\mathcal{L}_{\text{SFF}}$ is the semantic face feature loss and is responsible for the network to learn how to modify the semantic face features. It is defined as:
$\begin{split}&\mathcal{L}_{\text{SFF}}=\text{ MSE}(m_{j}^{\text{p}},m_{j}^{\text{t}})+\lambda_{\text{reg}}\ *\\ &\sum_{i=1,i\neq j}^{23}\text{MSE}(m_{i}^{\text{p}},m_{i}^{\text{t}})*(1-\lambda_{\text{cor}}\ *\ |c(m_{i},m_{j})|).\end{split}$ (3)
The first term calculates the MSE between the face feature $m_{j}^{p}$ of the edited face $I_{\text{edit}}$ and the target value for the feature $m_{j}^{t}$. To calculate $m_{j}^{p}$, the generated image is passed through a differentiable landmark detector in the forward-pass as illustrated in Figure 2 to predict the landmarks for $I_{\text{edit}}$. While this term encourages the model to change $w^{+}$ to manipulate the feature $m_{j}$, we must ensure that all other face features remain unchanged. Therefore, we add a regularising term with weighting $\lambda_{\text{reg}}$ that sums the MSE between predicted and target values for all other features where $m_{i}^{t}$ is fixed to the original image $I$. Furthermore, the regularising term is multiplied by the inverse of the absolute Person correlation coefficient $c(m_{i},m_{j})$ between the feature to be changed $m_{j}$ and those to be kept fixed $m_{i}$. Given that not all of our semantic features are necessarily fully disentangled given human anatomy, we estimate the correlations between the features on the FFHQ dataset [23] and introduce scalar $\lambda_{\text{cor}}$ that controls how much the correlation can relax the regularising term: If two features are strongly correlated, the regularising term between them is relaxed, and the network is less penalised for changing the correlated features $m_{i}$ together with the target feature $m_{j}$.